OpenStudy (adunb8):

this is differential equation & linear algebra level question. Exact Equation: Find the general solution, or find a relation that defines the solution implicitly. ex: (x^2-y)y' + 2x^3 +2xy = 0

OpenStudy (lgbasallote):

first step.. change y' into dy/dx \[\implies (x^2 - y)\frac {dy}{dx} + 2x^3 + 2xy = 0\] now multiply dx to all sides \[\implies (x^2 - y) dy + (2x^3 + 2xy)dx = 0\] following so far?

OpenStudy (adunb8):

yes

OpenStudy (lgbasallote):

now.. this is in teh form Mdx + Ndy = 0 and is exact... agree? the next step would be to integrate Mdx \[\implies \int Mdx\] in this case, our M is (2x^3 + 2xy) so our integral becomes \[\implies \int (2x^3 + 2xy)dx\] got it?

OpenStudy (adunb8):

yea i got it up to here.

OpenStudy (lgbasallote):

okay.. now if we simplify the integral...it becomes \[\frac{2x^4}{4} + \frac{2x^2 y}{2} + c\] so if you simplify that... \[\implies \frac{x^4}{2} + x^2 y + c\]

OpenStudy (lgbasallote):

now...the next step would be to replace c with g(y)<--g(y) is a function of y so we have.. \[\implies \frac{x^4}{2} + x^2 y + g(y)\]

OpenStudy (adunb8):

im following you! omg! then what happens?

OpenStudy (lgbasallote):

now... you derive that in terms of y so it becomes \[0 + x^2 + g^\prime(y)\] still okay?

OpenStudy (adunb8):

yes!!!

OpenStudy (lgbasallote):

so now you equate this to N in our case, N is (x^2 - y) <--because that's the function beside dy so we have \[\implies x^2 + g^\prime (y) = x^2 - y\] you can subtract both sides by x^2 that would give you.. \[g^\prime (y) = -y\] right?

OpenStudy (adunb8):

omg now i see! so what would be the final solution?

OpenStudy (lgbasallote):

you integrate both sides \[\implies \int g^\prime (y) = \int -y\] simplify it.. \[g(y) = -\frac{y^2}{2} + c\] now... remember the earlier expression? \[\frac{x^4}{2} + x^2 y + g(y)\] you substitute g(y) = -y^2/2 + c here so you have \[\implies \frac{x^4}{2} + x^2 y - \frac{y^2}{2} + c\] got it?

OpenStudy (adunb8):

so that would be the final solution? because the solution says something like \[y=x^2\pm \sqrt{2x^4+c}\]

OpenStudy (lgbasallote):

hmm seems this answer expressed it in a different way. let me try finding how to get it

OpenStudy (adunb8):

thank you so much =)

OpenStudy (lgbasallote):

where did you get that solution?

OpenStudy (adunb8):

its actually my teachers solution he made it himself. that is the reason i am so confused? i mean i would say its a typo if its little bit off but it looks like a completely different answer...

OpenStudy (lgbasallote):

okay i got it..

OpenStudy (adunb8):

this is the hw =)

OpenStudy (lgbasallote):

\[\frac{x^4}{2} + x^2y - \frac{y^2}{2} + c = 0\] multiply all terms by 2 \[\implies x^4 + 2x^2y - y^2 + c = 0\] transpose 2x^2y - y^2 \[\implies x^4 + c = y^2 - 2x^2 y\] add x^4 to both sides \[\implies x^4 + c + x^4 = y^2 - 2x^2 y + x^4\] combine like terms \[\implies 2x^4 + c = y^2 - 2x^2 y + x^4\] facotr out.. \[\implies 2x^4 + c = (y - x^2)^2\] take the square root of both sides \[\implies \pm \sqrt{2x^4 + c } + y - x^2\] add x^2 to both sides \[\implies y = x^2 \pm \sqrt{2x^4 + c}\] that's some filthy manipulation lol

OpenStudy (lgbasallote):

that second to the last line should be \[\implies \pm \sqrt{2x^4 + c} = y - x^2\] minor typo

OpenStudy (adunb8):

wow thats crazy... why would he do this to me .... =(

OpenStudy (lgbasallote):

calculus teachers are sadistic

OpenStudy (lgbasallote):

anyway...do you get the solution?

OpenStudy (adunb8):

basically my teacher wants to solve y= to something so i guess he wanted it ALL the way

OpenStudy (adunb8):

but yeah thank you so much you are my life saver without you i would automatically fail this course =)

OpenStudy (lgbasallote):

lol welcome. it's a great review for me too

OpenStudy (anonymous):

I guess my way sort of less painful: y² = x^4 + 2x²y + c -> y² - 2x²y = x^4 + c => ( y - x² ) ² = 2x^4 + c => y - x² = ± √ ( 2x^4 + c ) Thus y = x² ± √ ( 2x^4 + c )

OpenStudy (lgbasallote):

wasn't that what i did too?

OpenStudy (anonymous):

Might be the detailed explanation makes it looks long :P

OpenStudy (lgbasallote):

oh lol. i'll try to shorten it next time

OpenStudy (anonymous):

@lgbasallote No, you're always so cool at explanation. It's great benefit for the asker!

OpenStudy (adunb8):

i really like your explanation +)